OpenStudy (anonymous):

A circle is represented by the equation shown below: (x + 4)2 + (y + 6)2 = 9 Which statement is true? The circle is centered at (4, 6) and has a diameter of 9. The circle is centered at (-4, 6) and has a radius of 3. The circle is centered at (-4, -6) and has a radius of 3. The circle is centered at (4, -6) and has a diameter of 3.

jimthompson5910 (jim_thompson5910):

hint: the equation of a circle in general is \[\Large (x-h)^2+(y-k)^2=r^2\] this circle has the center (h,k) and the radius r

OpenStudy (anonymous):

so...

OpenStudy (skullpatrol):

Use ^ for exponentiation, like this (x + 4)^2 + (y + 6)^ 2 = 9, not just (x + 4)2 + (y + 6)2 = 9.

jimthompson5910 (jim_thompson5910):

For instance \[\Large (x-9)^2+(y-14)^2=25\] is in the form \[\Large (x-h)^2+(y-k)^2=r^2\] where (h,k) = (9,14) and r^2 = 25 ---> r = 5

OpenStudy (anonymous):

skull is that your real face?

OpenStudy (skullpatrol):

no

OpenStudy (anonymous):

o i was like DAMN!

OpenStudy (anonymous):

weird.........................................

OpenStudy (anonymous):

anyways haha

OpenStudy (anonymous):

so im a bit confused..

OpenStudy (anonymous):

lol your face on that little box is blurred by the sun :P

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

thanks?

OpenStudy (anonymous):

atleast i can tell your smiling :3

OpenStudy (anonymous):

OMG!!!!

OpenStudy (anonymous):

?.?

OpenStudy (anonymous):

I NEED HELP

OpenStudy (anonymous):

The circle is centered at (-4, -6) and has a radius of 3.

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

thats ur answer right there

OpenStudy (anonymous):

@aballer101 THANKS!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

ur welcome!

OpenStudy (anonymous):

NOW THAT HE POINTED OUT MY FACE IS BLUR IMA CHANGE IT :(

OpenStudy (anonymous):

sorry